# Shannon Capacity of Channels with Markov Insertions, Deletions and Substitutions

### v3

@@ V Capacity theorem for Markov-IDS channel
V-A Markov chain state distribution convergence
It is well-known (Theorem 8.9 in [21]), that if a Markov chain with a finite state space is aperiodic and irreducible, then
$\exists B\geq 0,0\leq P<1:\forall n\in{\mathbb{N}},\forall i,j\in{\mathcal{S}}:|G^{n}_{i,j}-\pi_{j}|\leq BP^{n}=B^{\prime}e^{-Cn}\stackrel{{\scriptstyle\triangle}}{{=}}\delta_{n},$ (31)
where $G^{n}_{i,j}=(G^{n})_{i,j}$, $\pi\in{\mathbb{D}}_{s}$ is the stationary distribution of Markov chain $G$: $\pi\cdot G=\pi$, constants $B,B^{\prime}\geq 0$, $10\leq P<$ 1 and $C>0$ depend only on the Markov chain.
For any starting distribution $\rho\in{\mathbb{D}}_{s}$ we have $\sum_{i}\rho_{i}=1$, and so
@@ V-A Markov chain state distribution convergence
Introduce sequence $\varepsilon_{n}$, such that
$\delta_{n}=\varepsilon_{n}\min_{i}\pi_{i}.$ (33)
Then, the small (with $n\to\infty$) additive term can be substituted with small multiplicative term as $\pi_{i}+\delta_{n}\leq\pi_{i}(1+\varepsilon_{n})$. Combining this with (31), (32)
$\forall\rho\in{\mathbb{D}}_{s},\forall j:~{}$ $\pi_{j}(1-\varepsilon_{n})\leq(\rho G^{n})_{j}\leq\pi_{j}(1+\varepsilon_{n}),$
@@
$\pi_{j}(1-\varepsilon_{n})\leq G^{n}_{i,j}\leq\pi_{j}(1+\varepsilon_{n}).$
V-B Main Theorems
@@
Proof.
We now will prove that the limit ${\mathbf{I}}(V[\rho])=\lim_{n\to\infty}{\mathbf{I}}(V_{n}[\rho])/n$ exists. For that, we will use the stronger formulation [22] of Fekete’s lemma, which says that if function $u:{\mathbb{N}}\to{\mathbb{R}}$ is quasi-subadditive in the sense that
$\forall m,n\in{\mathbb{N}},n\leq m\leq 2n:u(m+n)\leq u(m)+u(n)+f(m+n),$ (36)
where term $f(m+n)$ is such that series $\sum_{n=1}^{\infty}f(n)/n^{2}$ converges, then the limit $\lim_{n\to\infty}u(n)/n$ exists.
Denote by $V_{n_{1}|n_{2}|...|n_{t}}[\rho]:{\mathcal{X}}^{n}\rightsquigarrow{\overline{\mathcal{Y}}}^{t}$, $n=\sum_{i}n_{i}$, a channel, which works as a channel $V_{n,1}[\rho]$ combined with the merging operation. The merging operation concatenates the outputs inside $i$-th output sub-block for each $i$, where the corresponding $i$-th input sub-block has length $n_{i}$. So, the receiver knows the output, corresponding to each $n_{i}$-block, but does not know the exact output for each input symbol. Also, we write ${\widetilde{n}}_{i}$ instead of $n_{i}$ to denote the fact that the output, corresponding to the $i$-th input block is erased, i.e., substituted by the zero-length vector. Consider the case of $\rho=\pi$, i.e., the initial state distribution is the stationary distribution of the Markov chain (later we will show that the initial state distribution does not influence the i-capacity). Then, the following sequence of inequalities holds:
${\mathbf{I}}(V_{m+n}[\pi])$ $\stackrel{{\scriptstyle 1}}{{\leq}}{\mathbf{I}}(V_{m|n}[\pi])\stackrel{{\scriptstyle 2}}{{\leq}}{\mathbf{I}}(V_{m|l|n-l}[\pi])\stackrel{{\scriptstyle 3}}{{\leq}}{\mathbf{I}}(V_{m|{\widetilde{l}}|n-l}[\pi])+l{\overline{H}}\stackrel{{\scriptstyle 4}}{{\leq}}{\mathbf{I}}(V_{m|{\widetilde{l}}|n-l|l}[\pi])+l{\overline{H}}\stackrel{{\scriptstyle 5}}{{\leq}}{\mathbf{I}}(V_{m|{\widetilde{l}}|n}[\pi])+2l{\overline{H}}$ (37)
@@
In transition $3$, let us assume the same input distribution $p\in{\mathbb{D}}[{\mathcal{X}}^{m+n}]$ for both $V_{m|l|n-l}[\pi]$ and $V_{m|{\widetilde{l}}|n-l}[\pi]$ and prove that ${\mathbf{I}}(V_{m|l|n-l}[\pi],p)\leq{\mathbf{I}}(V_{m|{\widetilde{l}}|n-l}[\pi],p)+l{\overline{H}}$. Since this holds for any $p$, then we establish that the transition $3$ is also valid. Denote by $X_{1},X_{2},X_{3}$ the random variables corresponding to three input blocks of $m,l$ and $(n-l)$ symbols (they are the same for both channels since we take the same input distribution), and denote the random variables corresponding to the outputs by $Y_{1},Y_{2},Y_{3}$. Then, the amount of information about $X_{1}^{3}$ that we have lost by erasing $Y_{2}$ is given by
${\mathbf{I}}(V_{m|l|n-l}[\pi],p)-{\mathbf{I}}(V_{m|{\widetilde{l}}|n-l}[\pi],p)=I(Y_{1}^{3};X_{1}^{3})-I(Y_{1},Y_{3};X_{1}^{3})=I(Y_{2};X_{1}^{3}|Y_{1},Y_{3})\leq H(Y_{2}|Y_{1},Y_{3})\leq H(Y_{2}),$
@@
where $H(Y_{2})$ is the unconditional entropy of output corresponding to the second block; vector $Y_{2}$ consists of $l$ outputs $Y_{2}^{(i)}$ of component channels, glued together, $i=1\dots l$. So, $H(Y_{2})\leq H(Y_{2}^{(1)},Y_{2}^{(2)},\dots,Y_{2}^{(l)})\leq\sum_{i}H(Y_{2}^{(i)})$. Each $Y_{2}^{(i)}$ is an unconditioned output of a state channel. It can be expressed as
$H(Y_{2}^{(i)})={\mathbf{E}}_{\sigma_{i},x_{i}}{\left[{H(Y_{2}^{(i)}|x_{i},\sigma_{i})}\right]}\leq\max_{\sigma,x}{\mathbf{H}}{\left[{W[\sigma](y|x)}\right]}=\overline{H},$
@@
where distribution of $\sigma_{i}$ is given by initial state distribution $\rho$ and Markov chain transition probabilities, and distribution of $x_{i}$ is given by marginalization of distribution $p$ over all coordinates except the $(m+i)$-th coordinate. The above implies that $H(Y_{2})\leq l{\overline{H}}$ and inequality $3$ holds.
@@
In transition $4$ after block of $(n-l)$ symbols we further transmit a block of $l$ symbols. This cannot decrease the mutual information, since we can just ignore last block of input and output.
@@
In transition $5$ the output of $(n-l)$-block is merged with the output of the $l$-block. We can additionally tell the receiver the exact values of the output for the last $l$-block, and then the receiver can know the boundaries of the two outputs. By the same consideration as in inequality $3$, the conditional entropy about the last $l$ glued outputs, which we have partially lost by gluing, is not greater than unconditional entropy of $l$ unglued outputs, which is not greater than $l\overline{H}$ bits. This holds for any input distribution, so this holds for maxima over the input distributions.
@@
Now, consider channel $V_{m|{\widetilde{l}}|n}[\rho]:{\mathcal{X}}^{m+l+n}\rightsquigarrow{\overline{\mathcal{Y}}}^{2}$. We will need auxiliary channel $W_{n,1}[\sigma\to\tau]:{\mathcal{X}}^{n}\to{\mathcal{S}}\times{\overline{\mathcal{Y}}}^{n}$, where $\tau$ is considered as a part of the channel’s output. The channel can be represented as channel $W_{n,1}[\sigma]$ which additionally outputs the state after transmission of the last symbol:
$W_{n,1}[\sigma_{1}\shortrightarrow\tau](y_{1}^{n}|x_{1}^{n})$ $=W_{1}[\sigma_{1}](y_{1}|x_{1})\cdot\sum_{\sigma_{2}^{n}\in{\mathcal{S}}^{n-1}}G_{\sigma_{n},\tau}\cdot\prod_{i=2}^{n}G_{\sigma_{i-1},\sigma_{i}}W_{1}[\sigma](y_{i}|x_{i}).$ (38)
@@
The channel which does not output the final state and the channel which outputs the final state are related via marginalization over the final state:
$W_{n}[\sigma]({\overline{y}}|x)=\sum_{\tau}W_{n}[\sigma\shortrightarrow\tau]({\overline{y}}|x),~{}W_{n,1}[\sigma]({\overline{y}}|x)=\sum_{\tau}W_{n,1}[\sigma\shortrightarrow\tau]({\overline{y}}|x).$
@@
$V_{m|{\widetilde{l}}|n}[\pi](y_{1}^{2}|x_{1}^{3})\stackrel{{\scriptstyle\triangle}}{{=}}\sum_{\begin{subarray}{c}z_{1}\in\mathbf{g}^{-1}_{m}(y_{1})\\ z_{2}\in\mathbf{g}^{-1}_{n}(y_{2})\end{subarray}}\sum_{\sigma_{1}^{3}\in{\mathcal{S}}^{3}}\pi_{\sigma_{1}}W_{m}[\sigma_{1}\shortrightarrow\sigma_{2}]{\left({z_{1}|x_{1}}\right)}G^{l}_{\sigma_{2},\sigma_{3}}W_{n}[\sigma_{3}](z_{2}|x_{3})$
$\stackrel{{\scriptstyle 1}}{{\leq}}\sum_{z_{1}^{2},\sigma_{1}^{3}}~{}\pi_{\sigma_{1}}W_{m}[\sigma_{1}\shortrightarrow\sigma_{2}]{\left({z_{1}|x_{1}}\right)}\pi_{\sigma_{3}}(1+\varepsilon_{l})W_{n}[\sigma_{3}](z_{2}|x_{3})$
$\stackrel{{\scriptstyle 2}}{{\leq}}(1+\varepsilon_{l})\sum_{\sigma_{1},z_{1}}\pi_{\sigma_{1}}W_{m}[\sigma_{1}]{\left({z_{1}|x_{1}}\right)}\sum_{\sigma_{3},z_{2}}\pi_{\sigma_{3}}W_{n}[\sigma_{3}](z_{2}|x_{3})=(1+\varepsilon_{l})V_{m}[\pi](y_{1}|x_{1})V_{n}[\pi](y_{2}|x_{3}).$ (40)
@@
For brevity, introduce some shortcuts: let $x=x_{1}^{3}$, $y=y_{1}^{2}$. Since $V_{m|{\widetilde{l}}|n}[\pi](y_{1}^{2}|x_{1}^{3})$ does not depend on $x_{2}$, consider shortcut $U(y_{1}^{2}|x_{1}^{2})=V_{m|{\widetilde{l}}|n}[\pi](y_{1}^{2}|x_{1},x^{\prime},x_{2})$, where $x^{\prime}\in{\mathcal{X}}^{l}$ is some fixed vector. Also, consider shortcut $U^{\prime}(y|x)=V_{m}[\pi](y_{1}|x_{1})V_{n}[\pi](y_{2}|x_{2})$. Denote marginalized distribution ${\widetilde{p}}(x_{1},x_{2})=\sum_{x^{\prime}}p(x_{1},x^{\prime},x_{2})$. Rewriting (40) in shortcuts, one obtains $U(y|x)\leq(1+\varepsilon_{l})U^{\prime}(y|x)$, and one can upper-bound the mutual information by
${\mathbf{I}}(V_{m|{\widetilde{l}}|n}[\pi],p)={\mathbf{I}}(U,{\widetilde{p}})=\sum_{x,y}{\widetilde{p}}\circ U(x,y){\left[{\log_{2}U(y|x)-\log_{2}{\widetilde{p}}\circ U(*,y)}\right]}$
@@
$\stackrel{{\scriptstyle 1}}{{\leq}}\sum_{x,y}(1-\varepsilon_{l}){\widetilde{p}}\!\circ\!U^{\prime}(x,y)\log_{2}{\left[{(1+\varepsilon_{l})U^{\prime}(y|x)}\right]}-(1+\varepsilon_{l}){\widetilde{p}}\circ U^{\prime}(x,y)\log_{2}{\left[{(1-\varepsilon_{l}){\widetilde{p}}\!\circ\!U^{\prime}(*,y)}\right]}$
@@
$=(1-\varepsilon_{l}){\mathbf{E}}_{{\widetilde{p}}\circ U^{\prime}(x,y)}{\left[{\log_{2}U^{\prime}(y|x)+\log_{2}{\left({1+\varepsilon_{l}}\right)}}\right]}-(1+\varepsilon_{l}){\mathbf{E}}_{{\widetilde{p}}\circ U^{\prime}(x,y)}{\left[{\log_{2}{\widetilde{p}}\circ U^{\prime}(*,y)+\log_{2}{\left({1-\varepsilon_{l}}\right)}}\right]}$
@@
$={\mathbf{I}}(U^{\prime},{\widetilde{p}})-\varepsilon_{l}{\mathbf{E}}_{{\widetilde{p}}\circ U^{\prime}(x,y)}{\left[{\log_{2}U^{\prime}(y|x)}\right]}-\varepsilon_{l}{\mathbf{E}}{\left[{\log_{2}{\widetilde{p}}\circ U^{\prime}(*,y)}\right]}+\Gamma_{l}$
$={\mathbf{I}}(U^{\prime},{\widetilde{p}})+\varepsilon_{l}{\mathbf{H}}_{{\widetilde{p}}\circ U^{\prime}(x,y)}{\left[{U^{\prime}(y|x)}\right]}+\varepsilon_{l}{\mathbf{H}}{\left[{{\widetilde{p}}\circ U^{\prime}(*,y)}\right]}+\Gamma_{l}$
$\leq{\mathbf{I}}(V_{m}[\pi]\cdot V_{n}[\pi],{\widetilde{p}})+\varepsilon_{l}(m+n)\overline{H}+\varepsilon_{l}(m+n)\overline{H}+\Gamma_{l},$ (41)
@@
where $\Gamma_{l}=(1-\varepsilon_{l})\log_{2}(1+\varepsilon_{l})-(1+\varepsilon_{l})\log_{2}(1-\varepsilon_{l})$. The transition in $\stackrel{{\scriptstyle 1}}{{\leq}}$ is done as follows. Each $\log$ in the first term adds a negative amount, so the multiplicative term before the first logarithm is lower-bounded. The logarithm itself is an increasing function, so its argument is upper-bounded. Note that the upper bound might turn the value of logarithm to a positive value. Then the whole term under $\sum$ sign is upper-bounded, because the original term was non-positive. The second term is taken with a negative sign, so all bounds are the opposite.
Bringing together (37) and (41), for $u_{n}={\mathbf{I}}(V_{n}[\pi])$ one obtains
$u_{m+n}\leq u_{m}+u_{n}+2\varepsilon_{l}(m+n)\overline{H}+\Gamma_{l}+2l\overline{H}$
@@
$0\leq f(n)\leq 2\varepsilon_{2\sqrt{2n}}\cdot n+\Gamma_{\lfloor{{\sqrt{2n}}\rfloor}}+2\sqrt{2n}\cdot\overline{H}$
@@
Note that by (34), the first term tends to zero. The second term $\Gamma_{l}\to 0$ with $l\to\infty$. Thus, $f(n)\in O(\sqrt{n})$, so the series $\sum_{n}f(n)/n^{2}$ converges. This implies the existence of $\lim_{n\to\infty}u_{n}/n={\mathbf{I}}(V[\pi])$.
@@
We can show that ${\mathbf{I}}(V[\rho])={\mathbf{I}}(V[\pi])$ by the same trick of skipping $\lfloor{{\sqrt{n}}\rfloor}$ symbols at the beginning. ∎

### v5

@@ III-C Markov Chain State Distribution Convergence
It is well-known (Th. 8.9 in [32]), that if a Markov chain with a finite state space is aperiodic and irreducible, then there exist constants $B,B^{\prime}\geq 0$, $10\leq P<$ 1 and $C>0$, such that
$\forall n\in{\mathbb{N}},\forall i,j\in{\mathcal{S}}:|G^{n}_{i,j}-\pi_{j}|\leq BP^{n}=B^{\prime}e^{-Cn}{\overset{\triangle}{=}}\delta_{n},$ (32)
where $G^{n}_{i,j}=(G^{n})_{i,j}$, and $\pi\!\in\!{\mathbb{D}}_{s}$ is the stationary distribution of the Markov chain. Moreover, for any initial distribution $\rho$,
@@
Introduce sequence $\varepsilon_{n}=\delta_{n}/\min_{i}\pi_{i}$. Then, the small additive deviation (with $n\to\infty$) can be replaced with a small multiplicative deviation as $\pi_{i}+\delta_{n}\leq\pi_{i}(1+\varepsilon_{n})$. Combining this with (32) and (33),
$\forall\rho\in{\mathbb{D}}_{s},\forall j:~{}$ $\pi_{j}(1-\varepsilon_{n})\leq(\rho G^{n})_{j}\leq\pi_{j}(1+\varepsilon_{n}),$ (34)
@@
$\pi_{j}(1-\varepsilon_{n})\leq G^{n}_{i,j}\leq\pi_{j}(1+\varepsilon_{n}).$ (36)
III-D Main Theorems
@@
Assume that the length of the output of all state channels $W[\sigma]$ is constrained, i.e., $W[\sigma]:{\mathcal{X}}\rightsquigarrow{\widetilde{{\mathcal{Y}}}}$, where
${\widetilde{{\mathcal{Y}}}}=\cup_{i=0}^{A}{\mathcal{Y}}^{i},$ (38)
and $A$ is a constant which does not depend on $\sigma$.
Then, for any $\rho$, the channel sequence $V[\rho]$ is information stable. In particular, the coding capacity ${\mathbf{C}}(V[\rho])$ achieves the mutual information capacity ${\mathbf{I}}(V)$, and thus does not depend on $\rho$:
${\mathbf{C}}(V[\rho])={\mathbf{C}}(V)={\mathbf{I}}(V).$ (39)
The proofs of these theorems are given in the next section, employing several general results provided in the appendices.
In our development in Appendix A, we also introduce a unified approach to elucidate the idea behind Dobrushin’s methods, which can be seen as applying functions to channel sequences. If a function is applied to a channel’s output, and this function is not “doing much” in a certain sense, then the modified channel has the same mutual information capacity, coding capacity, and information stability as the original one. This means that one can apply any finite number of such functions and still have the same capacities.
In Appendix B, we provide two results that are also used in proving our theorems: Proposition 8 provides how perturbation of the sequence of optimal input distributions influences the asymptotic behavior of mutual information capacity. Proposition 9 gives the order of growth of the mutual information variance.
IV Proofs of the Main Theorems
IV-A Proof of Theorem 1
We first will prove that the limit ${\mathbf{I}}(V[\rho])$ exists. For that, we will use the stronger formulation [33] of Fekete’s lemma, which says that if function $u:{\mathbb{N}}\to{\mathbb{R}}$ is quasi-subadditive, i.e., if for all intergers $m,n$, such that $n\leq m\leq 2n$,
$u(m+n)\leq u(m)+u(n)+f(m+n),$ (40)
where function $f(n)$ is such that the series $\sum_{n=1}^{\infty}f(n)/n^{2}$ converges, then the limit $\lim_{n\to\infty}u(n)/n$ exists.
Denote by $V_{n_{1}|n_{2}|...|n_{t}}[\rho]:{\mathcal{X}}^{n}\rightsquigarrow{\overline{\mathcal{Y}}}^{t}$, $n=\sum_{i}n_{i}$, a channel, which works as a channel $V_{n,1}[\rho]$ combined with the merging operation. The merging operation concatenates the outputs inside $i$-th output sub-block for each $i$, where the corresponding $i$-th input sub-block has length $n_{i}$. Therefore, the receiver knows the output corresponding to each $n_{i}$ block but does not know the exact output for each input symbol. Also, we write ${\widetilde{n}}_{i}$ instead of $n_{i}$ to denote the fact that the output, corresponding to the $i$-th input block, is erased, i.e., substituted by the zero-length vector. Consider the case of $\rho=\pi$, i.e., the initial state distribution is the stationary distribution of the Markov chain (later, it will become clear that the initial state distribution does not influence the i-capacity). Then, the following sequence of inequalities holds:
${\mathbf{I}}(V_{m+n}[\pi])\stackrel{{\scriptstyle 1}}{{\leq}}{\mathbf{I}}(V_{m|n}[\pi])\stackrel{{\scriptstyle 2}}{{\leq}}{\mathbf{I}}(V_{m|l|n-l}[\pi])$
$\stackrel{{\scriptstyle 3}}{{\leq}}{\mathbf{I}}(V_{m|{\widetilde{l}}|n-l}[\pi])+l{\overline{H}}\stackrel{{\scriptstyle 4}}{{\leq}}{\mathbf{I}}(V_{m|{\widetilde{l}}|n-l|l}[\pi])+l{\overline{H}}$
$\stackrel{{\scriptstyle 5}}{{\leq}}{\mathbf{I}}(V_{m|{\widetilde{l}}|n}[\pi])+2l{\overline{H}}$ (41)
@@ IV-A Proof of Theorem 1
In transition $3$, assume the same input distribution $p\in{\mathbb{D}}[{\mathcal{X}}^{m+n}]$ for both $V_{m|l|n-l}[\pi]$ and $V_{m|{\widetilde{l}}|n-l}[\pi]$. We establish that the transition $3$ is valid by showing that ${\mathbf{I}}(V_{m|l|n-l}[\pi],p)\leq{\mathbf{I}}(V_{m|{\widetilde{l}}|n-l}[\pi],p)+l{\overline{H}}$ for any $p$. Denote by $X_{1},X_{2},X_{3}$ the random variables corresponding to three input blocks of $m,l$ and $(n-l)$ symbols (they are the same for both channels since we take the same input distribution), and denote the random variables corresponding to the outputs by $Y_{1},Y_{2},Y_{3}$. Then, the amount of information about $X_{1}^{3}$ that we have lost by erasing $Y_{2}$ is given by
${\mathbf{I}}(V_{m|l|n-l}[\pi],p)-{\mathbf{I}}(V_{m|{\widetilde{l}}|n-l}[\pi],p)$
$=I(Y_{1}^{3};X)-I(Y_{1},Y_{3};X)=I(Y_{2};X|Y_{1},Y_{3})\leq H(Y_{2}),$ (42)
@@
where $H(Y_{2})$ is the unconditional entropy of output corresponding to the second block. Vector $Y_{2}$ consists of $l$ outputs $Y_{2}^{(i)}$, $i=1\dots l$, of component channels, glued together, therefore $H(Y_{2})\leq H(Y_{2}^{(1)},Y_{2}^{(2)},\dots,Y_{2}^{(l)})\leq\sum_{i}H(Y_{2}^{(i)})$. Each entropy $H(Y_{2}^{(i)})$ is the unconditioned entropy of the output of a state channel. It can be expressed as
$H(Y_{2}^{(i)})\!=\!{\mathbf{E}}_{\sigma_{i},x_{i}}{\left[{H(Y_{2}^{(i)}|x_{i},\sigma_{i})}\right]}\leq\overline{H},$ (43)
@@
where distribution of $\sigma_{i}$ is given by initial state distribution $\rho=\pi$ and Markov chain transition probabilities, and distribution of $x_{i}$ is given by marginalization of distribution $p$ over all coordinates except the $(m+i)$-th coordinate. The above implies that $H(Y_{2})\leq l{\overline{H}}$ and inequality $3$ holds.
@@
In transition $4$, after a block of $(n-l)$ symbols, we transmit a block of $l$ symbols. This does not decrease mutual information.
@@
In transition $5$, the output of the $(n-l)$-block is merged with the output of the $l$-block. We can additionally tell the receiver the exact values of the output for the last $l$-block, and then the receiver can know the boundaries of the two outputs. By the same consideration as in inequality $3$, the conditional entropy about the last $l$ glued outputs, which we have partially lost by gluing, is not greater than the unconditional entropy of the $l$ unglued outputs, which is not greater than $l\overline{H}$ bits. This holds for any input distribution, so this holds for maxima over the input distributions.
@@
Now, consider channel $V_{m|{\widetilde{l}}|n}[\rho]:{\mathcal{X}}^{m+l+n}\rightsquigarrow{\overline{\mathcal{Y}}}^{2}$. We will need the auxiliary channel $W_{n,1}[\sigma\to\tau]:{\mathcal{X}}^{n}\to{\mathcal{S}}\times{\overline{\mathcal{Y}}}^{n}$, where $\tau$ is considered as a part of the channel’s output. The channel can be represented as channel $W_{n,1}[\sigma]$, which additionally outputs the state after transmission of the last symbol:
$W_{n,1}[\sigma_{1}\shortrightarrow\tau](y_{1}^{n}|x_{1}^{n})=W_{1}[\sigma_{1}](y_{1}|x_{1})$
×$\times\sum_{\sigma_{2}^{n}\in{\mathcal{S}}^{n-1}}G_{\sigma_{n},\tau}\cdot\prod_{i=2}^{n}G_{\sigma_{i-1},\sigma_{i}}W_{1}[\sigma](y_{i}|x_{i}).$ (44)
@@
The channel that does not output the final state and the channel that outputs the final state are related via marginalization over the final state:
$W_{n}[\sigma]({\overline{y}}|x)$ $=\sum_{\tau}W_{n}[\sigma\shortrightarrow\tau]({\overline{y}}|x),$ (46)
$W_{n,1}[\sigma]({\overline{y}}|x)$ $=\sum_{\tau}W_{n,1}[\sigma\shortrightarrow\tau]({\overline{y}}|x).$ (47)
@@
$\sum_{\begin{subarray}{c}z_{1}\in\mathbf{g}^{-1}_{m}(y_{1})\\ z_{2}\in\mathbf{g}^{-1}_{n}(y_{2})\end{subarray}}\sum_{\sigma_{1}^{3}\in{\mathcal{S}}^{3}}\!\!\pi_{\sigma_{1}}\!W_{m}[\sigma_{1}\!\shortrightarrow\!\sigma_{2}]{\left({z_{1}|x_{1}}\right)}G^{l}_{\sigma_{2},\sigma_{3}}W_{n}[\sigma_{3}](z_{2}|x_{3})$
$\leq\sum_{z_{1}^{2},\sigma_{1}^{3}}~{}\pi_{\sigma_{1}}W_{m}[\sigma_{1}\shortrightarrow\sigma_{2}]{\left({z_{1}|x_{1}}\right)}\pi_{\sigma_{3}}(1+\varepsilon_{l})W_{n}[\sigma_{3}](z_{2}|x_{3})$
$\leq(1+\varepsilon_{l})\sum_{\sigma_{1},z_{1}}\pi_{\sigma_{1}}W_{m}[\sigma_{1}]{\left({z_{1}|x_{1}}\right)}\sum_{\sigma_{3},z_{2}}\pi_{\sigma_{3}}W_{n}[\sigma_{3}](z_{2}|x_{3})$
$=(1+\varepsilon_{l})V_{m}[\pi](y_{1}|x_{1})V_{n}[\pi](y_{2}|x_{3}).$ (48)
@@
For brevity, introduce some shortcuts. Let $U(y_{1}^{2}|x_{1}^{2})=V_{m|{\widetilde{l}}|n}[\pi](y_{1}^{2}|x_{1},x^{\prime},x_{2})$. Since $V_{m|{\widetilde{l}}|n}[\pi](y_{1}^{2}|x_{1}^{3})$ does not depend on $x_{2}$, $x^{\prime}$ can be arbitrary. Also, let $U^{\prime}(y_{1}^{2}|x_{1}^{2})=V_{m}[\pi](y_{1}|x_{1})V_{n}[\pi](y_{2}|x_{2})$. Denote marginalized distribution ${\widetilde{p}}(x_{1},x_{2})=\sum_{x^{\prime}}p(x_{1},x^{\prime},x_{2})$.
Rewriting (48) in shortcuts, one obtains $U(y|x)\leq(1+\varepsilon_{l})U^{\prime}(y|x)$, and one can upper-bound the mutual information by (49) (at the bottom of this page), where $\Gamma_{l}=(1-\varepsilon_{l})\log_{2}(1+\varepsilon_{l})-(1+\varepsilon_{l})\log_{2}(1-\varepsilon_{l})$.
${\mathbf{I}}(V_{m|{\widetilde{l}}|n}[\pi],p)={\mathbf{I}}(U,{\widetilde{p}})=\sum_{x,y}{\widetilde{p}}\!\circ\!U(x,y){\left[{\log_{2}U(y|x)-\log_{2}{\widetilde{p}}\!\bullet\!U(y)}\right]}$
@@
$\stackrel{{\scriptstyle 6}}{{\leq}}\sum_{x,y}{\left\{{(1-\varepsilon_{l}){\widetilde{p}}\!\circ\!U^{\prime}(x,y)\log_{2}{\left[{(1+\varepsilon_{l})U^{\prime}(y|x)}\right]}-(1+\varepsilon_{l}){\widetilde{p}}\!\circ\!U^{\prime}(x,y)\log_{2}{\left[{(1-\varepsilon_{l}){\widetilde{p}}\!\bullet\!U^{\prime}(y)}\right]}}\right\}}$
@@
$=(1-\varepsilon_{l}){\mathbf{E}}_{{\widetilde{p}}\circ U^{\prime}(x,y)}{\left[{\log_{2}U^{\prime}(y|x)+\log_{2}{\left({1+\varepsilon_{l}}\right)}}\right]}-(1+\varepsilon_{l}){\mathbf{E}}_{{\widetilde{p}}\circ U^{\prime}(x,y)}{\left[{\log_{2}{\widetilde{p}}\!\bullet\!U^{\prime}(y)+\log_{2}{\left({1-\varepsilon_{l}}\right)}}\right]}$
@@
$={\mathbf{I}}(U^{\prime},{\widetilde{p}})-\varepsilon_{l}{\mathbf{E}}_{{\widetilde{p}}\circ U^{\prime}(x,y)}{\left[{\log_{2}U^{\prime}(y|x)}\right]}-\varepsilon_{l}{\mathbf{E}}{\left[{\log_{2}{\widetilde{p}}\!\bullet\!U^{\prime}(y)}\right]}+\Gamma_{l}={\mathbf{I}}(U^{\prime},{\widetilde{p}})+\varepsilon_{l}{\mathbf{H}}_{{\widetilde{p}}\circ U^{\prime}(x,y)}{\left[{U^{\prime}(y|x)}\right]}+\varepsilon_{l}{\mathbf{H}}{\left[{{\widetilde{p}}\!\bullet\!U^{\prime}(y)}\right]}+\Gamma_{l}$
$\leq{\mathbf{I}}(V_{m}[\pi]\cdot V_{n}[\pi],{\widetilde{p}})+\varepsilon_{l}(m+n)\overline{H}+\varepsilon_{l}(m+n)\overline{H}+\Gamma_{l}$ (49)
@@
The transition in $\stackrel{{\scriptstyle 6}}{{\leq}}$ is done as follows. Each $\log$ in the first term adds a negative amount, so the multiplicative term before the first logarithm is lower bounded. The logarithm itself is an increasing function, so its argument is upper-bounded. Note that the upper bound might turn the value of the logarithm to positive, hence lower bounding of the multiplicative term decreases the value of the summand. But since the original summand under $\sum$ was non-positive and became positive, it is still upper-bounded. The second term is taken with a negative sign, so all the bounds are the opposite.
Bringing together (41) and (49), for $u_{n}={\mathbf{I}}(V_{n}[\pi])$ we have
$u_{m+n}\leq u_{m}+u_{n}+2\varepsilon_{l}(m+n)\overline{H}+\Gamma_{l}+2l\overline{H}$ (50)
@@
$0\leq f(n)\leq 2\varepsilon_{2\sqrt{2n}}\cdot n+\Gamma_{\lfloor{{\sqrt{2n}}\rfloor}}+2\sqrt{2n}\cdot\overline{H}$ (51)
@@
Note that by (35), the first summand goes to zero. The second summand $\Gamma_{l}\to 0$ with $l\to\infty$. Thus, $f(n)\in O(\sqrt{n})$, so the series $\sum_{n}f(n)/n^{2}$ converges. This implies the existence of $\lim_{n\to\infty}u_{n}/n={\mathbf{I}}(V[\pi])$.
@@
We can show that ${\mathbf{I}}(V[\rho])={\mathbf{I}}(V[\pi])$ by the same trick of skipping $\lfloor{{\sqrt{n}}\rfloor}$ symbols at the beginning.